OpenStudy (teller):

Can someone make sure I did this right? A certain forest covers an area of 4300km^2. Suppose that each year this area decreases by 7.75%. What will the area be after 15 years?Use the calculator provided and round your answer to the nearest square kilometer. p=4300km^2 r=0.075 t=15 F=430KM^2(1-0.0775)^15 = 128.22245212km2

OpenStudy (triciaal):

typo 430 s/b 4300 I think right formula

OpenStudy (teller):

they gave us the formula of f=p(1+r)^t or f=p(1-r)^t

OpenStudy (teller):

from what I know I plugged it in right.

OpenStudy (triciaal):

I did not verify the calculation

OpenStudy (phi):

tricaal is saying use 4300km^2. You used 430

OpenStudy (teller):

Oh sorry

OpenStudy (teller):

1282.22452118km2 ?

OpenStudy (phi):

round your answer to the nearest square kilometer.

OpenStudy (teller):

so would it be 1283?

OpenStudy (triciaal):

no to round when less than 1/2 way to what you need you drop

OpenStudy (teller):

Huh?

OpenStudy (phi):

is 1282.2 closer to 1282 or 1283 ?

OpenStudy (teller):

1282

OpenStudy (triciaal):

here you have 1282.2 and you need whole number 1282 then 1283

OpenStudy (teller):

so 1283?

OpenStudy (phi):

the rule is if you have 1282.5 (5 or bigger in the 10ths place) round up otherwise, just drop the decimal yes 1282 is the rounded answer.

OpenStudy (teller):

alright thank you

OpenStudy (triciaal):